# Positive definiteness and Fell bundles over discrete groups

### v1

@@ 3.3 Categorical considerations
The category ${\tt Fell}$-${\tt PD}$ contains a subcategory ${\tt Fell}$ with the same objects, but where an arrow $(\varphi,T)$ from ${\mathcal{A}}$ to ${\mathcal{B}}$ is required to be a morphism according to Definition 3.1. As $\Phi_{T}=\mathcal{F}(\varphi,T)$ is a *-homomorphism from $C^{*}({\mathcal{A}})$ into $C^{*}({\mathcal{B}})$ whenever $(\varphi,T)$ is a morphism, we may restrict $\mathcal{F}$ to ${\tt Fell}$ and get a functor from ${\tt Fell}$ into the category ${\tt C^{*}HOM}$ of $C^{*}$-algebras having *-homomorphisms as arrows.
We may also introduce the subcategory ${\tt Fell}$-${\tt PD}_{\tt red}$ of ${\tt Fell}$-${\tt PD}$ having the same objects but whose arrows involve only group homomorphisms with amenable kernels. The fact that composition of arrows in this category makes sense is due to following result: if $\varphi:G\to H$ and $\psi:H\to K$ are group homomorphisms such that ${\rm ker}(\varphi)$ and ${\rm ker}(\psi)$ are amenable, then ${\rm ker}(\psi\varphi)$ is amenable. This can be shown by first observing that ${\rm ker}(\varphi)$ is a normal subgroup of ${\rm ker}(\psi\varphi)$ such that the quotient group ${\rm ker}(\psi\varphi)/{\rm ker}(\varphi)$ is isomorphic to a subgroup of ${\rm ker}(\psi)$, from which the amenability of ${\rm ker}(\psi\varphi)$ then follows by using that amenability is preserved when passing to subgroups and closed under extensions.
We then get a functor $\mathcal{F}_{\rm red}$ from ${\tt Fell}$-${\tt PD}_{\tt red}$ to ${\tt C^{*}CP}$ by setting
$\mathcal{F}_{\rm red}({\mathcal{A}}):=C_{r}^{*}({\mathcal{A}})\text{ and }\mathcal{F}_{\rm red}(\varphi,T):=M_{T}:C_{r}^{*}({\mathcal{A}})\to C_{r}^{*}({\mathcal{B}})$
for an object ${\mathcal{A}}$ in ${\tt Fell}$-${\tt PD}_{\tt red}$ and an arrow $(\varphi,T)$ from ${\mathcal{A}}$ to ${\mathcal{B}}$ in ${\tt Fell}$-${\tt PD}_{\tt red}$.
The category ${\tt Fell}$-${\tt PD}_{\tt red}$ contains an obvious subcategory ${\tt Fell}_{\tt red}$ having the same objects, but whose arrows are morphisms $(\varphi,T):{\mathcal{A}}\to{\mathcal{B}}$ with ${\rm ker}(\varphi)$ is amenable. By restricting $\mathcal{F}_{\rm red}$ to ${\tt Fell}_{\tt red}$ we get a functor from ${\tt Fell}_{\tt red}$ into ${\tt C^{*}HOM}$.
Sometimes, the subcategory ${\tt C^{*}CCP}$ of ${\tt C^{*}CP}$ whose arrows consist of contractive completely positive maps is the one of interest. It follows from Theorems 3.10 and 3.12 that we obtain functors into this category by restricting $\mathcal{F}$ and $\mathcal{F}_{\rm red}$ to the respective subcategories of ${\tt Fell}$-${\tt PD}$ and ${\tt Fell}$-${\tt PD}_{\tt red}$ whose arrows are such that $T_{e}$ is contractive.
@@ 4 Some examples
Let’s first consider the simple case where ${\mathcal{A}}=({\mathbb{C}}\times\{g\})_{g\in G}$ is the group bundle associated to $G$, =×${\mathcal{B}}={\mathbb{C}}\times\{e\}$ is the trivial bundle over $\{e\}$ and $\varphi_{0}:G\to\{e\}$ is the trivial homomorphism. Then one readily sees that an ${\mathcal{A}}$-$\varphi_{0}$$-{\mathcal{B}}$ bundle map $T=(T_{g})_{g\in G}$ is of the form $T_{g}(z,g)=(\phi(g)z,e)$ for some (necessarily bounded) function $\phi:G\to{\mathbb{C}}$, and that $T$ is positive definite if and only if $\phi$ is positive definite. In this case, we have $C^{*}({\mathcal{A}})=C^{*}(G)$, $C^{*}({\mathcal{B}})={\mathbb{C}}$, and the map $T\mapsto\Phi_{T}$ corresponds to the usual correspondence between positive definite functions on $G$ and positive linear functionals on $C^{*}(G)$.
Moreover, we have $C_{r}^{*}({\mathcal{A}})=C_{r}^{*}(G)$, $C_{r}^{*}({\mathcal{B}})={\mathbb{C}}$. Thinking of positive definite ${\mathcal{A}}$-$\varphi_{0}$$-{\mathcal{B}}$ bundle maps as positive definite functions on $G$, it is clear that $M_{T}$ is defined exactly when $T$ is a positive definite function associated with a unitary representation of $G$ weakly contained in $\lambda^{G}$, in which case $M_{T}$ is the corresponding positive linear functional on $C_{r}^{*}(G)$. Note that this is true regardless of whether $G={\rm ker}(\varphi_{0})$ is amenable or not.
@@
We now look at a more general situation, covering both the previous examples. Let ${\mathcal{A}}=(A_{g})_{g\in G}$ be a Fell bundle and $B$ be a $C^{*}$-algebra. Let ${\mathcal{B}}$ be the Fell bundle over the trivial group $\{e\}$ with fiber $B$ over $e$, so $B=C^{*}({\mathcal{B}})$, and let $\varphi_{0}:G\to\{e\}$ be the trivial homomorphism. Then a positive definite ${\mathcal{A}}$-$\varphi_{0}$$-{\mathcal{B}}$ bundle map $T=(T_{g})_{g\in G}$ is a family of linear maps $T_{g}:A_{g}\to B$ such that the matrix $\big{[}T_{g_{i}^{-1}g_{j}}(a_{i}^{*}a_{j})\big{]}$ is positive in $M_{n}(B)$ for every $g_{1},\ldots,g_{n}\in G$ and $a_{i}\in A_{g_{i}},i=1,\ldots,n$.
In particular, choosing $B={\mathbb{C}}$, we get that a positive definite ${\mathcal{A}}$-$\varphi_{0}$$-{\mathbb{C}}$ bundle map $\omega=(\omega_{g})_{g\in G}$ is a family of linear functionals $\omega_{g}:A_{g}\to{\mathbb{C}}$ such that the matrix $\big{[}\omega_{g_{i}^{-1}g_{j}}(a_{i}^{*}a_{j})\big{]}$ is positive in $M_{n}({\mathbb{C}})$ for every $g_{1},\ldots,g_{n}\in G$ and $a_{i}\in A_{g_{i}},i=1,\ldots,n$. This is easily seen to be equivalent to requiring that
@@
for every $g_{1},\ldots,g_{n}\in G$ and $a_{i}\in A_{g_{i}},i=1,\ldots,n$. Such a bundle map $\omega=(\omega_{g})_{g\in G}$ is essentially the same as what Fell and Doran call a functional on ${\mathcal{A}}$ of positive type, cf. [29, Chapter VIII, 20.2]. We refer to [4] for a recent study of such maps. In accordance with Theorem 3.10, every such a functional $\omega=(\omega_{g})_{g\in G}$ on ${\mathcal{A}}$ gives rise to a positive linear functional $\Phi_{\omega}$ on $C^{*}({\mathcal{A}})$ satisfying that $\Phi_{\omega}(\,\widehat{j}_{g}^{\mathcal{A}}(a))=\omega_{g}(a)$ for $g\in G$ and $a\in A_{g}$. Conversely, each positive linear functional $\Psi$ on $C^{*}({\mathcal{A}})$ gives rise to a positive definite ${\mathcal{A}}$-$\varphi_{0}$$-{\mathbb{C}}$ bundle map $\omega^{\Psi}=(\omega^{\Psi}_{g})_{g\in G}$ given by $\omega_{g}^{\Psi}(a)=\Psi(\,\widehat{j}_{g}^{\mathcal{A}}(a))$ for $g\in G$ and $a\in A_{g}$, which satisfies that $\Psi=\Phi_{\omega^{\Psi}}$. It follows that the map $\omega\mapsto\Phi_{\omega}$ is a bijection between the cone of positive definite ${\mathcal{A}}$-$\varphi_{0}$$-{\mathbb{C}}$ bundle maps and the cone of positive linear functionals on $C^{*}({\mathcal{A}})$.
In another direction, let now ${\mathcal{A}}=({\mathbb{C}}\times\{g\})_{g\in G}$ be the group bundle over $G$ and assume $B=B({\mathcal{H}})$ for some Hilbert space ${\mathcal{H}}$. Then a positive definite ${\mathcal{A}}$-$\varphi_{0}$$-{\mathcal{B}}$ bundle map amounts to a map $T:G\to B({\mathcal{H}})$ which is completely positive definite in the sense of Paulsen [56, Chapter 4, p. 51] (some other authors just say positive definite, e.g. [13, Section 6, p. 115] or [53, Section 8.3]), i.e., it satisfies that the matrix $[T(g_{i}^{-1}g_{j})]$ is positive in $M_{n}(B({\mathcal{H}}))$ for every $g_{1},\ldots,g_{n}\in G$. As mentioned by Paulsen on p. 54, there is a one-to-one correspondence between such completely positive definite maps from $G$ into $B({\mathcal{H}})$ and completely positive maps from $C^{*}(G)$ into $B({\mathcal{H}})$. This fits with Theorem 3.10.
Example 4.4.
We can now conclude from Theorem 3.10 that there is a unique completely positive linear map $\Phi_{T}:C^{*}({\mathcal{A}})\to C^{*}({\mathcal{B}})$ satisfying that
$\Phi_{T}(\widehat{j}^{\,{\mathcal{A}}}_{g}(a))=\widehat{j}^{\,{\mathcal{B}}}_{\varphi(g)}(T_{g}(a))$
for all $g\in G$ and $a\in A_{g}$. As $T$ is multiplicative, one readily checks that $\phi_{T}:C_{c}({\mathcal{A}})\to C_{c}({\mathcal{B}})$ is multiplicative. Since its extension $\Phi_{T}$ is bounded, this implies that $\Phi_{T}$ is multiplicative, hence that $\Phi_{T}$ is a *-homomorphism.
Similarly, if one assumes that ${\rm ker}(\varphi)$ is amenable, we get from Theorem 3.12 that there is a unique *-homomorphism $M_{T}:C_{r}^{*}({\mathcal{A}})\to C_{r}^{*}({\mathcal{B}})$ satisfying that
$M_{T}(\lambda^{\mathcal{A}}_{g}(a))=\lambda^{\mathcal{B}}_{g}(T_{g}(a))$
for all $g\in G$ and $a\in A_{g}$.
These observations generalize [27, Propositions 21.2 and 21.3], where the case $G=H$ and $\varphi={\rm id}_{G}$ is considered.
@@
where $a_{ij}:=c_{i}b_{i}^{*}b_{j}c_{j}^{*}\in B_{e}$. Now, the matrix $[\phi(g_{i}^{-1}g_{j})]$ is positive in $M_{n}({\mathbb{C}})$ if $\phi$ is positive definite, and $[a_{ij}]$ is clearly positive in $M_{n}(B_{e})$. Hence, the matrix $[\phi(g_{i}^{-1}g_{j})a_{ij}]$ is then positive in $M_{n}(B_{e})$ (cf. [47, Lemma 3.1]) and the claim readily follows.
Thus, Theorems 3.10 and 3.12 apply and give completely positive linear maps $\Phi^{\phi}:C^{*}({\mathcal{B}})\to C^{*}({\mathcal{B}})$ and $M^{\phi}:C_{r}^{*}({\mathcal{B}})\to C_{r}^{*}({\mathcal{B}})$ determined by
@@
Note that if $\phi:G\to{\mathbb{T}}$ is a character, then $T^{\phi}$ is easily seen to be a morphism from ${\mathcal{B}}$ into itself, and the associated map $\Phi^{\gamma}$ (resp. $M^{\gamma}$) is a *-automorphism of $C^{*}({\mathcal{B}})$ (resp. $C_{r}^{*}({\mathcal{B}})$).
@@
Let again ${\mathcal{B}}=(B_{g})_{g\in G}$ be Fell bundle. We may then also consider maps of the type used by Exel to introduce the approximation property for Fell bundles [26, 27], cf. Example 5.2. Let $\xi:G\to B_{e}$ be a finitely supported function. For each $g\in G$, define a linear map $T_{g}:B_{g}\to B_{g}$ by
@@
Then $T=(T_{g})_{g\in G}$ is a positive definite ${\mathcal{B}}$-bundle map.
Indeed, let $g_{1},\ldots,g_{n}\in G$ and $b_{i},c_{i}\in B_{g_{i}}$ for $i=1,\ldots,n$. For each $j\in\{1,\ldots,n\}$, let $d_{j}:G\to B_{e}$ be the finitely supported function defined by
@@
$(a,g)(a^{\prime},g^{\prime}):=(a\alpha_{g}(a^{\prime}),gg^{\prime}),$
and
$(a,g)^{*}:=(\alpha_{g^{-1}}(a^{*}),g^{-1}),$
respectively. Each $B_{h}=\{(b,h):b\in B\}$ is defined similarly, along with the Fell bundle operations.
@@
for all $a\in A$. Then $T=(T_{g})_{g\in G}$ is easily seen to be an ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map. Moreover, $T$ is positive definite if and only if the family $(\widetilde{T}_{g})_{g\in G}$ is $\Sigma$-$\varphi$-$\Omega$ positive definite in the following sense: for every $g_{1},\ldots,g_{n}\in G$ and $a_{1},\ldots,a_{n}\in A$, the matrix
@@
is positive in $M_{n}(B)$. [^1] When $G=H$, $\varphi={\rm id}_{G}$, $\Sigma=\Omega$ and $A$ is unital, this amounts to say that $(\tilde{T}_{g})_{g\in G}$ is positive definite w.r.t. $\Sigma$ in the sense of [9].
Indeed, by definition, $T=(T_{g})_{g\in G}$ is positive definite if and only if
$\sum_{i,j=1}^{n}(b_{i},\varphi(g_{i}))\,T_{g_{i}^{-1}g_{j}}\big{(}(a_{i},g_{i})^{*}(a_{j},g_{j})\big{)}\,(b_{j},\varphi(g_{j}))^{*}\,\in\,B^{+}\times\{e\}$
whenever $g_{1},\ldots,g_{n}\in G$, $a_{1},\ldots,a_{n}\in A$ and $b_{1},\ldots,b_{n}\in B$. On the other hand, $(\widetilde{T}_{g})_{g\in G}$ is $\Sigma$-$\varphi$-$\Omega$ positive definite if and only if
$\sum_{i,j=1}^{n}b_{i}\,\beta_{\varphi(g_{i})}\big{(}\widetilde{T}_{g_{i}^{-1}g_{j}}\big{(}\alpha_{{g_{i}}^{-1}}({a_{i}}^{*}a_{j})\big{)}\big{)}\,{b_{j}}^{*}\,\in B^{+}$
whenever $g_{1},\ldots,g_{n}\in G$ and $a_{1},\ldots,a_{n}\in A$ and $b_{1},\ldots,b_{n}\in B$. As we have
$~(b_{i},\varphi(g_{i}))$ $\,T_{g_{i}^{-1}g_{j}}\big{(}(a_{i},g_{i})^{*}(a_{j},g_{j})\big{)}\,(b_{j},\varphi(g_{j}))^{*}$
$=(b_{i},\varphi(g_{i}))\,T_{g_{i}^{-1}g_{j}}\big{(}(\alpha_{{g_{i}}^{-1}}({a_{i}}^{*}),{g_{i}}^{-1})(a_{j},g_{j})\big{)}\,(\beta_{{\varphi(g_{j})}^{-1}}({b_{j}}^{*}),\varphi({g_{j}})^{-1})$
$=(b_{i},\varphi(g_{i}))\,T_{g_{i}^{-1}g_{j}}\big{(}(\alpha_{{g_{i}}^{-1}}({a_{i}}^{*})\alpha_{{g_{i}}^{-1}}(a_{j}),{g_{i}}^{-1}g_{j})\big{)}\,(\beta_{{\varphi(g_{j})}^{-1}}({b_{j}}^{*}),\varphi({g_{j}})^{-1})$
$=(b_{i},\varphi(g_{i}))\,T_{g_{i}^{-1}g_{j}}\big{(}(\alpha_{{g_{i}}^{-1}}({a_{i}}^{*}a_{j}),{g_{i}}^{-1}g_{j})\big{)}\,(\beta_{{\varphi(g_{j})}^{-1}}({b_{j}}^{*}),\varphi({g_{j}})^{-1})$
$=(b_{i},\varphi(g_{i}))\,\big{(}\widetilde{T}_{g_{i}^{-1}g_{j}}\big{(}\alpha_{{g_{i}}^{-1}}({a_{i}}^{*}a_{j})\big{)},\varphi({g_{i}}^{-1}g_{j})\big{)}\,(\beta_{{\varphi(g_{j})}^{-1}}({b_{j}}^{*}),\varphi({g_{j}})^{-1})$
$=\big{(}b_{i}\,\beta_{\varphi(g_{i})}\big{(}\widetilde{T}_{g_{i}^{-1}g_{j}}\big{(}\alpha_{{g_{i}}^{-1}}({a_{i}}^{*}a_{j})\big{)}\big{)},\varphi(g_{j})\big{)}\,(\beta_{{\varphi(g_{j})}^{-1}}({b_{j}}^{*}),\varphi({g_{j}})^{-1})$
$=\big{(}b_{i}\,\beta_{\varphi(g_{i})}\big{(}\widetilde{T}_{g_{i}^{-1}g_{j}}\big{(}\alpha_{{g_{i}}^{-1}}({a_{i}}^{*}a_{j})\big{)}\big{)}\,{b_{j}}^{*},e\big{)},$
the claim readily follows.
Applying Theorem 3.10, we get that $(\widetilde{T}_{g})_{g\in G}$ is $\Sigma$-$\varphi$-$\Omega$ positive definite if and only if there exists a completely positive map $\Phi_{\widetilde{T}}:C^{*}(\Sigma)\to C^{*}(\Omega)$ such that
@@
Moreover, if we assume that ${\rm ker}(\varphi)$ is amenable, then Theorem 3.12 gives that $(\widetilde{T}_{g})_{g\in G}$ is $\Sigma$-$\varphi$-$\Omega$ positive definite if and only if there exists a completely positive map $M_{\widetilde{T}}:C_{r}^{*}(\Sigma)\to C_{r}^{*}(\Omega)$ such that
$M_{\widetilde{T}}(a\lambda_{\Sigma}(g))=\widetilde{T}_{g}(a)\lambda_{\Omega}(\varphi(g))$
for all $a\in A$ and $g\in G$, in which case $\|M_{\widetilde{T}}\|=\|\widetilde{T}_{e}\|$.
We note that in the case where $G=H$, $\varphi={\rm id}_{G}$, $\Sigma=\Omega$ and $A$ is unital, these two results boil down to the equivalence of (b), (c) and (d) in [9, Corollary 4.4]. Under the same assumptions, we also illustrated in [9, Example 4.1] how positive definite families $(\widetilde{T}_{g})_{g\in G}$ may be constructed from equivariant representations of $\Sigma$. We will now describe how such families arise when $\Omega$ is possibly different from $\Sigma$ and $A$ may be non-unital, still assuming that $G=H$ and $\varphi={\rm id}_{G}$.
@@
for each $g\in G,a\in A,b\in B$ and $x,y\in X$. (This means that $\gamma$ is an $\alpha$-$\beta$ compatible action of $G$ on the right Hilbert $A$-$B$-bimodule $X$ in the sense of [24], except that we don’t assume non-degeneracy of the left action of $A$ on $X$.)
For each $g\in G$ define a linear map $\widetilde{T}_{g}:A\to B$ by
$\widetilde{T}_{g}(a)=\big{\langle}x,a\cdot(\gamma_{g}(x))\big{\rangle}_{B}$
for all $a\in A$. Then $(\widetilde{T}_{g})_{g\in G}$ is $\Sigma$-${\rm id}_{G}$-$\Omega$ positive-definite. Indeed, let $g_{1},\ldots,g_{n}\in G$ and $a_{1},\ldots,a_{n}\in A$. Then for $i,j\in\{1,\ldots,n\}$ we have
$\beta_{g_{i}}\Big{(}\widetilde{T}_{g_{i}^{-1}g_{j}}\big{(}\alpha_{{g_{i}}^{-1}}(a_{i}^{*}a_{j})\big{)}\Big{)}$ $=\beta_{g_{i}}\Big{(}\big{\langle}x,\alpha_{{g_{i}}^{-1}}(a_{i}^{*}a_{j})\cdot(\gamma_{g_{i}^{-1}g_{j}}(x))\big{\rangle}_{B}\Big{)}$
$=\big{\langle}\gamma_{g_{i}}(x),\gamma_{g_{i}}\big{(}\alpha_{{g_{i}}^{-1}}(a_{i}^{*}a_{j})\cdot\gamma_{g_{i}^{-1}g_{j}}(x)\big{)}\big{\rangle}_{B}$
$=\big{\langle}\gamma_{g_{i}}(x),\alpha_{g_{i}}(\alpha_{{g_{i}}^{-1}}(a_{i}^{*}a_{j}))\cdot\gamma_{g_{i}}(\gamma_{g_{i}^{-1}g_{j}}(x))\big{\rangle}_{B}$
$=\big{\langle}\gamma_{g_{i}}(x),(a_{i}^{*}a_{j})\cdot\gamma_{g_{j}}(x)\big{\rangle}_{B}$
$=\big{\langle}a_{i}\cdot\gamma_{g_{i}}(x),a_{j}\cdot\gamma_{g_{j}}(x)\big{\rangle}_{B}$
$=\big{\langle}y_{i},y_{j}\rangle_{B}$
where $y_{i}:=a_{i}\cdot\gamma_{g_{i}}(x)\in X$ for every $i$. Thus,
$\Big{[}\beta_{g_{i}}\big{(}\widetilde{T}_{g_{i}^{-1}g_{j}}(\alpha_{{g_{i}}^{-1}}(a_{i}^{*}a_{j}))\big{)}\Big{]}=\big{[}\big{\langle}y_{i},y_{j}\rangle_{B}\big{]}\,\in M_{n}(B)^{+},$
as desired.
We note that by adapting the proof of [9, Theorem 4.5], one can show that that every $\Sigma$-${\rm id}_{G}$-$\Omega$ positive definite family may be obtained in such a way, at least when $A$ and $B$ are unital.

### v2

@@ 3.2 Completely positive maps
Moreover, ${\tt Fell}$-${\tt PD}$ contains a subcategory ${\tt Fell}$ with the same objects, but where arrows are required to be morphisms according to Definition 3.1. As $\Phi_{T}=\mathcal{F}(\varphi,T)$ is then a *-homomorphism from $C^{*}({\mathcal{A}})$ into $C^{*}({\mathcal{B}})$, we may restrict $\mathcal{F}$ to ${\tt Fell}$ and get a functor into the category ${\tt C^{*}HOM}$ of $C^{*}$-algebras having *-homomorphisms as arrows.
Similarly, we may also consider the subcategory ${\tt Fell}$-${\tt PD}_{\tt red}$ of ${\tt Fell}$-${\tt PD}$ having the same objects but whose arrows involve only group homomorphisms with amenable kernels, and obtain a functor $\mathcal{F}_{\rm red}$ from ${\tt Fell}$-${\tt PD}_{\tt red}$ to ${\tt C^{*}CP}$. By restricting $\mathcal{F}_{\rm red}$ to the subcategory ${\tt Fell}_{\tt red}$ having the same objects, but whose arrows are morphisms $(\varphi,T):{\mathcal{A}}\to{\mathcal{B}}$ with ${\rm ker}(\varphi)$ is amenable, we get a functor from ${\tt Fell}_{\tt red}$ into ${\tt C^{*}HOM}$.
@@ 4 Some examples
We first consider the simple case where ${\mathcal{A}}=({\mathbb{C}}\times\{g\})_{g\in G}$ is the group bundle associated to $G$, =×${\mathcal{B}}={\mathbb{C}}\times\{e\}$ is the trivial bundle over $\{e\}$ and $\varphi_{0}:G\to\{e\}$ is the trivial homomorphism. Then one readily sees that an ${\mathcal{A}}$-$\varphi_{0}$$-{\mathcal{B}}$ bundle map $T=(T_{g})_{g\in G}$ is of the form $T_{g}(z,g)=(\phi(g)z,e)$ for some (necessarily bounded) function $\phi:G\to{\mathbb{C}}$, and that $T$ is positive definite if and only if $\phi$ is positive definite. In this case, we have $C^{*}({\mathcal{A}})=C^{*}(G)$, $C^{*}({\mathcal{B}})={\mathbb{C}}$, and the map $T\mapsto\Phi_{T}$ corresponds to the usual correspondence between positive definite functions on $G$ and positive linear functionals on $C^{*}(G)$.
Moreover, we have $C_{r}^{*}({\mathcal{A}})=C_{r}^{*}(G)$, $C_{r}^{*}({\mathcal{B}})={\mathbb{C}}$. Thinking of a positive definite ${\mathcal{A}}$-$\varphi_{0}$$-{\mathcal{B}}$ bundle map $T$ as a positive definite function $\phi$ on $G$, it follows for example from [37, Lemma 2.1.2] that $M_{T}$ exists as a positive linear functional on $C_{r}^{*}(G)$ exactly when the unitary representation of $G$ associated with $\phi$ via the GNS (or Gelfand-Raikov) construction is weakly contained in $\lambda^{G}$ (in the sense of [22, 37]). Note that this is true regardless of whether $G={\rm ker}(\varphi_{0})$ is amenable or not.
@@
We now look at a more general situation, covering both previous examples. Let ${\mathcal{A}}=(A_{g})_{g\in G}$ be a Fell bundle and $B$ be a $C^{*}$-algebra. Let ${\mathcal{B}}$ be the Fell bundle over the trivial group $\{e\}$ with fiber $B$ over $e$, so $B=C^{*}({\mathcal{B}})$, and let $\varphi_{0}:G\to\{e\}$ be the trivial homomorphism. Then a positive definite ${\mathcal{A}}$-$\varphi_{0}$$-{\mathcal{B}}$ bundle map $T=(T_{g})_{g\in G}$ is a family of linear maps $T_{g}:A_{g}\to B$ such that the matrix $\big{[}T_{g_{i}^{-1}g_{j}}(a_{i}^{*}a_{j})\big{]}$ is positive in $M_{n}(B)$ for every $g_{1},\ldots,g_{n}\in G$ and $a_{i}\in A_{g_{i}},i=1,\ldots,n$.
In particular, choosing $B={\mathbb{C}}$, we get that a positive definite ${\mathcal{A}}$-$\varphi_{0}$$-{\mathbb{C}}$ bundle map $\omega=(\omega_{g})_{g\in G}$ is a family of bounded linear functionals $\omega_{g}:A_{g}\to{\mathbb{C}}$ such that the matrix $\big{[}\omega_{g_{i}^{-1}g_{j}}(a_{i}^{*}a_{j})\big{]}$ is positive in $M_{n}({\mathbb{C}})$ for every $g_{1},\ldots,g_{n}\in G$ and $a_{i}\in A_{g_{i}},i=1,\ldots,n$. This is easily seen to be equivalent to requiring that
@@
for every $g_{1},\ldots,g_{n}\in G$ and $a_{i}\in A_{g_{i}},i=1,\ldots,n$. Such a bundle map $\omega=(\omega_{g})_{g\in G}$ is essentially the same as what Fell and Doran call a functional on ${\mathcal{A}}$ of positive type, cf. [30, Chapter VIII, 20.2]. We refer to [4] for a recent study of such maps. In accordance with Theorem 3.11, every such a functional $\omega=(\omega_{g})_{g\in G}$ on ${\mathcal{A}}$ gives rise to a positive linear functional $\Phi_{\omega}$ on $C^{*}({\mathcal{A}})$ satisfying that $\Phi_{\omega}(\,\widehat{j}_{g}^{\mathcal{A}}(a))=\omega_{g}(a)$ for $g\in G$ and $a\in A_{g}$. Conversely, each positive linear functional $\Psi$ on $C^{*}({\mathcal{A}})$ gives rise to a positive definite ${\mathcal{A}}$-$\varphi_{0}$$-{\mathbb{C}}$ bundle map $\omega^{\Psi}=(\omega^{\Psi}_{g})_{g\in G}$ given by $\omega_{g}^{\Psi}(a)=\Psi(\,\widehat{j}_{g}^{\mathcal{A}}(a))$ for $g\in G$ and $a\in A_{g}$, which satisfies that $\Psi=\Phi_{\omega^{\Psi}}$. It follows that the map $\omega\mapsto\Phi_{\omega}$ is a bijection between the cone of positive definite ${\mathcal{A}}$-$\varphi_{0}$$-{\mathbb{C}}$ bundle maps and the cone of positive linear functionals on $C^{*}({\mathcal{A}})$.
In another direction, let now ${\mathcal{A}}=({\mathbb{C}}\times\{g\})_{g\in G}$ be the group bundle over $G$ and assume $B=B({\mathcal{H}})$ for some Hilbert space ${\mathcal{H}}$. Then a positive definite ${\mathcal{A}}$-$\varphi_{0}$$-{\mathcal{B}}$ bundle map amounts to a map $T:G\to B({\mathcal{H}})$ which is completely positive definite in the sense of Paulsen [57, Chapter 4, p. 51] (some other authors just say positive definite, e.g. [13, Section 6, p. 115] or [53, Section 8.3]), i.e., it satisfies that the matrix $[T(g_{i}^{-1}g_{j})]$ is positive in $M_{n}(B({\mathcal{H}}))$ for every $g_{1},\ldots,g_{n}\in G$. As mentioned by Paulsen on p. 54, there is a one-to-one correspondence between such completely positive definite maps from $G$ into $B({\mathcal{H}})$ and completely positive maps from $C^{*}(G)$ into $B({\mathcal{H}})$. This fits with Theorem 3.11.
@@
where $a_{ij}:=c_{i}b_{i}^{*}b_{j}c_{j}^{*}\in B_{e}$. Now, the matrix $[\phi(g_{i}^{-1}g_{j})]$ is positive in $M_{n}({\mathbb{C}})$ since $\phi$ is positive definite, and $[a_{ij}]$ is clearly positive in $M_{n}(B_{e})$. Hence, the matrix $[\phi(g_{i}^{-1}g_{j})a_{ij}]$ is then positive in $M_{n}(B_{e})$ (cf. [23, Lemma 3.1]) and the claim readily follows.
Thus, Theorems 3.11 and 3.14 apply and give completely positive linear maps $\Phi^{\phi}:C^{*}({\mathcal{B}})\to C^{*}({\mathcal{B}})$ and $M^{\phi}:C_{r}^{*}({\mathcal{B}})\to C_{r}^{*}({\mathcal{B}})$ determined by
@@
Note that if $\gamma:G\to{\mathbb{T}}$ is a character, then $T^{\gamma}$ is easily seen to be a morphism from ${\mathcal{B}}$ into itself, and the associated map $\Phi^{\gamma}$ (resp. $M^{\gamma}$) is a *-automorphism of $C^{*}({\mathcal{B}})$ (resp. $C_{r}^{*}({\mathcal{B}})$).
@@
Let ${\mathcal{B}}=(B_{g})_{g\in G}$ be Fell bundle. We may then consider maps of the type used by Exel to introduce the approximation property for Fell bundles [27, 28], see also Example 5.2. Let $\xi:G\to B_{e}$ be a finitely supported function. For each $g\in G$, define a linear map $T_{g}:B_{g}\to B_{g}$ by
@@
Then $T=(T_{g})_{g\in G}$ is a positive definite ${\mathcal{B}}$-bundle map. Indeed, let $g_{1},\ldots,g_{n}\in G$ and $b_{i},c_{i}\in B_{g_{i}}$ for $i=1,\ldots,n$. For each $j\in\{1,\ldots,n\}$, let $d_{j}:G\to B_{e}$ be the finitely supported function defined by
@@
[…]
respectively. Each fiber $B_{h}$ of ${\mathcal{B}}$ is defined similarly, along with the Fell bundle operations.
@@
for all $a\in A$. Then $T=(T_{g})_{g\in G}$ is easily seen to be an ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map. Moreover, a straightforward computation gives that $T$ is positive definite if and only if the family $(\widetilde{T}_{g})_{g\in G}$ is $\Sigma$-$\varphi$-$\Omega$ positive definite in the following sense: for every $g_{1},\ldots,g_{n}\in G$ and $a_{1},\ldots,a_{n}\in A$, the matrix
@@
is positive in $M_{n}(B)$. [^1] When $G=H$, $\varphi={\rm id}_{G}$, $\Sigma=\Omega$ and $A$ is unital, this amounts to say that $(\tilde{T}_{g})_{g\in G}$ is positive definite w.r.t. $\Sigma$ in the sense of [10]. Applying Theorem 3.11, we get that $(\widetilde{T}_{g})_{g\in G}$ is $\Sigma$-$\varphi$-$\Omega$ positive definite if and only if there exists a completely positive map $\Phi_{\widetilde{T}}:C^{*}(\Sigma)\to C^{*}(\Omega)$ such that
@@
If ${\rm ker}(\varphi)$ is amenable, then we may apply Theorem 3.14 and obtain a similar result in the reduced case. When $G=H$, $\varphi={\rm id}_{G}$, $\Sigma=\Omega$ and $A$ is unital, these two results boil down to the equivalence of (b), (c) and (d) in [10, Corollary 4.4]. Under the same assumptions, we also illustrated in [10, Example 4.1] how positive definite families $(\widetilde{T}_{g})_{g\in G}$ may be constructed from equivariant representations of $\Sigma$. We will now describe how such families arise when $\Omega$ is possibly different from $\Sigma$ and $A$ may be non-unital, still assuming that $G=H$ and $\varphi={\rm id}_{G}$.
@@
for all $g\in G,a\in A,b\in B$ and $x,y\in X$. (This means that $\gamma$ is an $\alpha$-$\beta$ compatible action of $G$ on the right Hilbert $A$-$B$-bimodule $X$ in the sense of [25], except that we do not assume non-degeneracy of the left action of $A$ on $X$.) For each $g\in G$ define a linear map $\widetilde{T}_{g}:A\to B$ by
$\widetilde{T}_{g}(a)=\big{\langle}x,a\cdot\gamma_{g}(x)\big{\rangle}_{B}$
for all $a\in A$. Then one easily verifies that $(\widetilde{T}_{g})_{g\in G}$ is $\Sigma$-${\rm id}_{G}$-$\Omega$ positive-definite. By adapting the proof of [10, Theorem 4.5], it can be shown that every $\Sigma$-${\rm id}_{G}$-$\Omega$ positive definite family may be obtained in such a way, at least when $A$ and $B$ are unital.